OpenStudy (anonymous):

Factor. Check by multiplying factors Medal rewarded 8x^3+2x^2+12x+3

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

@hartnn @Law&Order

hartnn (hartnn):

8x^3+2x^2 factor out 2x^2 from those 2 terms, what do u get ?

OpenStudy (anonymous):

4x^2

hartnn (hartnn):

what, how ? \(\Large 8x^2 +2x^2 = 2x^2\times 4x +2x^2 = 2x^2 (4x+1) \) got this ?

OpenStudy (anonymous):

Oh I see what you did there.

hartnn (hartnn):

now factor out 3 from 12x+3

OpenStudy (anonymous):

4x

hartnn (hartnn):

example of factoring : \(\large ☻☺+☻♥= ☻(☺+♥) \\ \text{this is factoring, } \\ ☻\text{ was common from both terms and factored out.}\)

hartnn (hartnn):

just 4x ?

hartnn (hartnn):

\(\Large 12x+3 =3\times 4x+3 = 3 (4x+1)\)

OpenStudy (anonymous):

OH +3. Okay.

hartnn (hartnn):

now you have \(\Large 2x^2 (4x+1)+3(4x+1)\) what do u get when you factor out 4x+1 ??

OpenStudy (anonymous):

(2x^2+3)(4x+1)

hartnn (hartnn):

\(\huge \color{red}{\checkmark \quad \ddot \smile}\)

hartnn (hartnn):

can you verify your answer by multiplying these factors? or need help with that too ?

hartnn (hartnn):

you should again get 8x^3+2x^2+12x+3 from (2x^2+3)(4x+1)

OpenStudy (anonymous):

2x^2(4x+1) 8x^2+2x^2

OpenStudy (anonymous):

Yeah all that lol I got it, Thanks alot man!

hartnn (hartnn):

welcome ^_^